# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ Abstract
We develop a family of mixed finite element methods for a model of nonlinear poroelasticity where, thanks to a rewriting of the constitutive equations, the permeability depends on the total poroelastic stress and on the fluid pressure and therefore we can use the Hellinger–Reissner principle with weakly imposed stress symmetry for Biot’s equations. The problem is adequately structured into a coupled system consisting of one saddle-point formulation, one linearised perturbed saddle-point formulation, and two off-diagonal perturbations. This system’s unique solvability requires assumptions on regularity and Lipschitz continuity of the inverse permeability, and the analysis follows fixed-point arguments and the Babuška–Brezzi theory. The discrete problem is shown uniquely solvable by applying similar fixed-point and saddle-point techniques as for the continuous case. The method is based on the classical PEERSk elements, it is exactly momentum and mass conservative, and it is robust with respect to the nearly incompressible as well as vanishing storativity limits. We derive a priori error estimates, we also propose fully computable residual-based a posteriori error indicators, and show that they are reliable and efficient with respect to the natural norms, and robust in the limit of near incompressibility. These a posteriori error estimates are used to drive adaptive mesh refinement. The theoretical analysis is supported and illustrated by several numerical examples in 2D and 3D.
@@ 1 Introduction
Our interest is in deriving mixed finite element (FE) formulations (solving also for other variables of interest), and for this we can cite in particular [36, 38], where fully mixed formulations based on the Hu–Washizu principle are studied. Writing the poroelasticity equations in terms of the strain tensor was motivated in particular in [38] because the permeability – at least in the regime we focus here – depends nonlinearly on the total amount of fluid, which is a function of strain.
The upshot here compared to [38] is that we are able to rewrite the constitutive equation for permeability to depend on the total poroelastic stress and on the fluid pressure (similarly as in, e.g., [8]). This allows us to revert to the more popular Hellinger–Reissner type of mixed formulations for poroelasticity [7, 39, 40, 14, 46] (without solving explicitly for the strain). Consequently, another appealing advantage with respect to the formulation in [38] is that, as in the Hellinger–Reissner formulation, the model becomes robust with respect to the Lamé constants. Also in contrast to [38], in this work we use a mixed form for the fluid flow (adding the discharge flux as additional unknown), which gives the additional advantage of mass conservativity.
@@
An additional goal of this work is to derive efficient and reliable residual a posteriori error estimators for the nonlinear poroelasticity equations. The approach follows a similar treatment as that of [32] (which focuses on mixed formulations of stress-assisted diffusion equations), with the difference that here we do not need to include augmentation terms for the mixed form of the mixed diffusion problem. The main ingredients in the analysis of these estimates are Helmholtz decompositions and a global inf-sup (together with boundedness and Lipschitz continuity of coupling terms), local inverse and trace estimates, bubble-based localisation arguments, and properties of Clément and Raviart–Thomas interpolators. See also [3, 28] for estimators in a similar multiphysics context and, e.g., [19, 20, 41] for mixed linear elasticity. Note that for the reliability of the estimator the aforementioned Helmholtz decompositions – for both tensor-vector and vector-scalar cases – should be valid for mixed boundary conditions. For this we follow [3] and [25], from which we inherit an a convexity assumption on the Neumann sub-boundary (where we impose traction and flux boundary conditions).
Outline. The rest of the paper is organised as follows. The remainder of this Section has a collection of preliminary definitions and notational convention, as well as the statement of the governing partial differential equations. The weak formulation and proofs of the uniform boundedness of the bilinear forms and suitable inf-sup conditions are shown in Section 2. The fixed-point analysis of the coupled problem is carried out in Section 3. Section 4 then focuses on the Galerkin discretisation, including its well-posedness analysis and definition of specific FE subspaces that provide momentum and mass conservativity. In Section 5 we show a Céa estimate and using appropriate approximation properties we derive optimal a priori error bounds including also the higher order case. The definition of a residual a posteriori error estimator and the proofs of its reliability and efficiency are presented in Section 6. We conclude in Section 7 with some numerical tests that both validate and underline the theoretical properties of the proposed discretisations.
@@
$\mathbb{L}_{\mathrm{skew}}^{2}(\Omega):=\{\boldsymbol{\tau}\in\mathbb{L}^{2}(\Omega):\,\boldsymbol{\tau}=-\boldsymbol{\tau}^{\tt t}\},$
@@
As usual $\mathbb{I}$ stands for the identity tensor in $\mathbb{R}^{d\times d}$, and ||$\cdot$|| ⋅ | denotes the Euclidean norm in $\mathbb{R}^{d}$. Also, for any vector field $\boldsymbol{v}=(v_{i})_{i=1,d}$ we set the gradient and divergence operators as $\boldsymbol{\nabla}\boldsymbol{v}\,:=\,\left(\frac{\partial v_{i}}{\partial x_{j}}\right)_{i,j=1,d}$ and $\mathop{\rm div}\nolimits\boldsymbol{v}\,:=\,\sum_{j=1}^{d}\frac{\partial v_{j}}{\partial x_{j}}$. In addition, for any tensor fields $\boldsymbol{\tau}=(\tau_{ij})_{i,j=1,d}$ and ${\boldsymbol{\zeta}}=(\zeta_{ij})_{i,j=1,d}$, we let $\mathbf{div}\,\boldsymbol{\tau}$ be the divergence operator $\mathop{\rm div}\nolimits$ acting along the rows of $\boldsymbol{\tau}$, and define the transpose, the trace, the tensor inner product, and the deviatoric tensor as $\boldsymbol{\tau}^{\mathrm{t}}:=(\tau_{ji})_{i,j=1,d}$, $\mathop{\rm tr}(\boldsymbol{\tau}):=\sum_{i=1}^{d}\tau_{ii}$, $\boldsymbol{\tau}:{\boldsymbol{\zeta}}:=\sum_{i,j=1}^{n}\tau_{ij}\zeta_{ij}$, and $\boldsymbol{\tau}^{\mathrm{d}}:=\boldsymbol{\tau}-\frac{1}{n}\,\mathop{\rm tr}(\boldsymbol{\tau})\,\mathbb{I}$, respectively. We also recall the Hilbert space
@@
$\mathbb{H}(\mathbf{div};\Omega)\,:=\,\big{\{}\boldsymbol{\tau}\in\mathbb{L}^{2}(\Omega):\,\mathbf{div}\,\boldsymbol{\tau}\in\mathbf{L}^{2}(\Omega)\big{\}}.$
@@
The balance of linear momentum for the solid-fluid mixture is written as
@@
with $\boldsymbol{\sigma}$ being the total Cauchy stress tensor of the mixture (sum of the effective solid and fluid stresses), whose dependence on strain and on fluid pressure is given by the constitutive assumption (or effective stress principle)
(1.2) $\boldsymbol{\sigma}=\mathcal{C}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})-\alpha p\mathbb{I}\qquad\text{in $\Omega$}.$
@@
We also consider the balance of angular momentum, which in this context states that the total poroelastic stress is a symmetric tensor $\boldsymbol{\sigma}=\boldsymbol{\sigma}^{\tt t}$. To weakly impose it, it is customary to use the rotation tensor
(1.3) $\boldsymbol{\rho}=\frac{1}{2}(\boldsymbol{\nabla}\boldsymbol{u}-[\boldsymbol{\nabla}\boldsymbol{u}]^{\tt t})=\boldsymbol{\nabla}\boldsymbol{u}-\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}).$
@@
(1.4) $\zeta=c_{0}p+\alpha\mathop{\rm div}\nolimits\boldsymbol{u},$
where $c_{0}\geq 0$ is the constrained specific storage coefficient. Using Darcy’s law to describe the discharge velocity in terms of the fluid pressure gradient, the balance of mass for the total amount of fluid is $\partial_{t}\zeta-\mathop{\rm div}\nolimits(\kappa\nabla p)=g$ in $\Omega\times(0,t_{\mathrm{end}})$, where $\kappa$ is the intrinsic permeability of the medium, a nonlinear function of the porosity. In turn, in the small strains limit the porosity can be approximated by a linear function of the fluid content $\zeta$ (see for example [45, Section 2.1]), and so, thanks to (1.4), we can simply write $\kappa(\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}),p)$. Furthermore, after a backward Euler semi-discretisation in time with a constant time step and rescaling appropriately, we only consider the type of equations needed to solve at each time step and therefore we will concentrate on the form
(1.5) $c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})-\mathop{\rm div}\nolimits(\kappa(\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}),p)\nabla p)=g\qquad\text{in $\Omega$}.$
@@
(1.6) $\kappa(\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}),p)=\frac{k_{0}}{\mu_{f}}\mathbb{I}+\frac{k_{1}}{\mu_{f}}\exp(k_{2}(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})))\mathbb{I},\quad\kappa(\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}),p)=\frac{k_{0}}{\mu_{f}}\mathbb{I}+\frac{k_{1}(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u}))^{3}}{\mu_{f}(1-(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})))^{2}}\mathbb{I},$
where $\mu_{f}$ denotes the viscosity of the interstitial fluid and $k_{0},k_{1},k_{2}$ are model constants. We note that in the case of incompressible constituents one has $c_{0}=0$ and $\alpha=1$, indicating that permeability depends only on the dilation $\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})=\mathop{\rm div}\nolimits\boldsymbol{u}$ (see, e.g., [11]). We also note that even in such a scenario (of incompressible phases) the overall mixture is not necessarily incompressible itself. More precise assumptions on the behaviour of the permeability are postponed to Section 2.2. Next we note that from (1.2) we can obtain
(1.7) $\mathop{\rm tr}\boldsymbol{\sigma}=(d\lambda+2\mu)\mathop{\rm div}\nolimits\boldsymbol{u}-d\alpha p{\quad\hbox{and}\quad}\mathcal{C}^{-1}\boldsymbol{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbb{I}=\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})\qquad\text{in $\Omega$}.$
@@
$\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}(\boldsymbol{u})=\dfrac{1}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}+\dfrac{d\alpha}{d\lambda+2\mu}p,$
@@
(1.8) $\begin{array}[]{cc}\displaystyle\kappa(\boldsymbol{\sigma},p)=\frac{k_{0}}{\mu_{f}}\mathbb{I}+\frac{k_{1}}{\mu_{f}}\exp\Big{(}\dfrac{k_{2}}{d\lambda+2\mu}\big{(}(c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\boldsymbol{\sigma}\big{)}\Big{)}\mathbb{I},\\[4.30554pt] \displaystyle\kappa(\boldsymbol{\sigma},p)=\frac{k_{0}}{\mu_{f}}\mathbb{I}+\frac{k_{1}\big{(}(c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\boldsymbol{\sigma}\big{)}^{3}}{(d\lambda+2\mu)\mu_{f}\big{(}d\lambda+2\mu-((c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\boldsymbol{\sigma})\big{)}^{2}}\mathbb{I}.\end{array}$
@@
(1.9) $\mathcal{C}^{-1}\boldsymbol{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbb{I}=\boldsymbol{\nabla}\boldsymbol{u}-\boldsymbol{\rho}\qquad\text{in $\Omega$}.$
Finally, we introduce the discharge flux ${\boldsymbol{\varphi}}$ as an unknown defined by the constitutive relation
(1.10) $\kappa(\boldsymbol{\sigma},p)^{-1}{\boldsymbol{\varphi}}=\nabla p,$
@@
(1.11) $c_{0}p+\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}p-\mathop{\rm div}\nolimits{\boldsymbol{\varphi}}=g\qquad\text{in $\Omega$}.$
To close the system, we consider mixed boundary conditions for a given $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1/2}(\Gamma_{\mathrm{D}})$ and $p_{\mathrm{D}}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{D}})$:
(1.12) $\boldsymbol{u}=\boldsymbol{u}_{\mathrm{D}}{\quad\hbox{and}\quad}p=p_{\mathrm{D}}{\quad\hbox{on}\quad}\Gamma_{\mathrm{D}},\qquad{\boldsymbol{\varphi}}\cdot\mbox{\boldmath{$n$}}=0{\quad\hbox{and}\quad}\boldsymbol{\sigma}\mbox{\boldmath{$n$}}=\boldsymbol{0}{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}.$

### v2

# Khan, Lamichhane, Ruiz-Baier, Villa-Fuentes

@@ Abstract
We develop a family of mixed finite element methods for a model of nonlinear poroelasticity where, thanks to a rewriting of the constitutive equations, the permeability depends on the total poroelastic stress and on the fluid pressure and therefore we can use the Hellinger–Reissner principle with weakly imposed stress symmetry for Biot’s equations. The problem is adequately structured into a coupled system consisting of one saddle-point formulation, one linearised perturbed saddle-point formulation, and two off-diagonal perturbations. This system’s unique solvability requires assumptions on regularity and Lipschitz continuity of the inverse permeability, and the analysis follows fixed-point arguments and the Babuška–Brezzi theory. The discrete problem is shown uniquely solvable by applying similar fixed-point and saddle-point techniques as for the continuous case. The method is based on the classical PEERSk elements, it is exactly equilibrium and mass conservative, and it is robust with respect to the nearly incompressible as well as vanishing storativity limits. We derive a priori error estimates, we also propose fully computable residual-based a posteriori error indicators, and show that they are reliable and efficient with respect to the natural norms, and robust in the limit of near incompressibility. These a posteriori error estimates are used to drive adaptive mesh refinement. The theoretical analysis is supported and illustrated by several numerical examples in 2D and 3D.
@@ 1 Introduction
Our interest is in deriving mixed finite element (FE) formulations (solving also for other variables of interest), and for this we can cite in particular [37, 39], where fully mixed formulations based on the Hu–Washizu principle are studied. Writing the poroelasticity equations in terms of the strain tensor was motivated in particular in [39] because the permeability – at least in the regime we focus on here – depends nonlinearly on the total amount of fluid, which is a function of strain.
The upshot here compared to [39] is that we are able to rewrite the constitutive equation for permeability to depend on the total poroelastic stress and on the fluid pressure (similarly as in, e.g., [8]). This allows us to revert to the more popular Hellinger–Reissner type of mixed formulations for poroelasticity [7, 40, 41, 14, 47] (without solving explicitly for the strain). Consequently, another appealing advantage with respect to the formulation in [39] is that, as in the Hellinger–Reissner formulation, the model becomes robust in the nearly incompressibility regime. Also in contrast to [39], in this work we use a mixed form for the fluid flow (adding the discharge flux as additional unknown), which gives the additional advantage of mass conservativity.
@@
An additional goal of this work is to derive efficient and reliable residual a posteriori error estimators for the nonlinear poroelasticity equations. The approach follows a similar treatment as that of [33] (which focuses on mixed formulations of stress-assisted diffusion equations), with the difference that here we do not need to include augmentation terms for the mixed form of the mixed diffusion problem. The main ingredients in the analysis of these estimates are Helmholtz decompositions and a global inf-sup (together with boundedness and Lipschitz continuity of coupling terms), local inverse and trace estimates, bubble-based localisation arguments, and properties of Clément and Raviart–Thomas interpolators. See also [3, 28] for estimators in a similar multiphysics context and, e.g., [19, 20, 42] for mixed linear elasticity. Note that for the reliability of the estimator the aforementioned Helmholtz decompositions – for both tensor-vector and vector-scalar cases – should be valid for mixed boundary conditions. For this we follow [3] and [25], from which we inherit a convexity assumption on the Neumann sub-boundary (where we impose traction and flux boundary conditions).
Outline. The rest of the paper is organised as follows. The remainder of this Section has a collection of preliminary definitions and notational convention, as well as the statement of the governing partial differential equations. The weak formulation and proofs of the uniform boundedness of the bilinear forms and suitable inf-sup conditions are shown in Section 2. The fixed-point analysis of the coupled problem is carried out in Section 3. Section 4 then focuses on the Galerkin discretisation, including its well-posedness analysis and definition of specific FE subspaces that provide equilibrium and mass conservativity. In Section 5 we show a Céa estimate and using appropriate approximation properties we derive optimal a priori error bounds including also the higher order case. The definition of a residual a posteriori error estimator and the proofs of its reliability and efficiency are presented in Section 6. We display in Section 7 some numerical tests that both validate and underline the theoretical properties of the proposed discretisations, and close in Section 8 with a summary and a discussion on possible extensions.
@@
$\mathbb{L}_{\mathrm{skew}}^{2}(\Omega):=\{\bm{\tau}\in\mathbb{L}^{2}(\Omega):\,\bm{\tau}=-\bm{\tau}^{\tt t}\},$
@@
As usual ${\mathbf{I}}$ stands for the identity tensor in $\mathbb{R}^{d\times d}$, and ||$\cdot$|| ⋅ | denotes the Euclidean norm in $\mathbb{R}^{d}$. Also, for any vector field $\mathbf{v}=(v_{i})_{i=1,d}$ we set the gradient and divergence operators as $\bm{\nabla}\mathbf{v}\,:=\,\left(\frac{\partial v_{i}}{\partial x_{j}}\right)_{i,j=1,d}$ and $\mathop{\rm div}\nolimits\mathbf{v}\,:=\,\sum_{j=1}^{d}\frac{\partial v_{j}}{\partial x_{j}}$. In addition, for any tensor fields $\bm{\tau}=(\tau_{ij})_{i,j=1,d}$ and ${\bm{\zeta}}=(\zeta_{ij})_{i,j=1,d}$, we let $\mathbf{div}\,\bm{\tau}$ be the divergence operator $\mathop{\rm div}\nolimits$ acting along the rows of $\bm{\tau}$, and define the transpose, the trace, the tensor inner product, and the deviatoric tensor as $\bm{\tau}^{\mathrm{t}}:=(\tau_{ji})_{i,j=1,d}$, $\mathop{\rm tr}(\bm{\tau}):=\sum_{i=1}^{d}\tau_{ii}$, $\bm{\tau}:{\bm{\zeta}}:=\sum_{i,j=1}^{n}\tau_{ij}\zeta_{ij}$, and $\bm{\tau}^{\mathrm{d}}:=\bm{\tau}-{\frac{1}{d}}\,\mathop{\rm tr}(\bm{\tau})\,\mathbf{I}$, respectively. We also recall the Hilbert space
@@
$\mathbb{H}(\mathbf{div};\Omega)\,:=\,\big{\{}\bm{\tau}\in\mathbb{L}^{2}(\Omega):\,\mathbf{div}\,\bm{\tau}\in\mathbf{L}^{2}(\Omega)\big{\}}.$
@@
The equilibrium (balance of linear momentum) for the solid-fluid mixture is written as
@@
with $\bm{\sigma}$ being the total Cauchy stress tensor of the mixture (sum of the effective solid and fluid stresses), whose dependence on strain and on fluid pressure is given by the constitutive assumption (or effective stress principle)
$\bm{\sigma}=\mathcal{C}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})-\alpha p{\mathbf{I}}\qquad\text{in $\Omega$}.$ (1.2)
@@
We also consider the balance of angular momentum, which in this context states that the total poroelastic stress is a symmetric tensor $\bm{\sigma}=\bm{\sigma}^{\tt t}$. To weakly impose it, it is customary to use the rotation tensor
$\bm{\rho}=\frac{1}{2}(\bm{\nabla}{\bm{u}}-[\bm{\nabla}{\bm{u}}]^{\tt t})=\bm{\nabla}{\bm{u}}-\mbox{\boldmath{$\varepsilon$}}({\bm{u}}).$ (1.3)
@@
$\zeta=c_{0}p+\alpha\mathop{\rm div}\nolimits{\bm{u}},$ (1.4)
where $c_{0}\geq 0$ is the constrained specific storage coefficient. Using Darcy’s law to describe the discharge velocity in terms of the fluid pressure gradient, the balance of mass for the total amount of fluid is $\partial_{t}\zeta-\mathop{\rm div}\nolimits({\bm{\kappa}}\nabla p)=g$ in $\Omega\times(0,t_{\mathrm{end}})$, where ${\bm{\kappa}}$ is the intrinsic permeability tensor of the medium, a nonlinear function of the porosity. In turn, in the small strains limit the porosity can be approximated by a linear function of the fluid content $\zeta$ (see for example [46, Section 2.1]), and so, thanks to (1.4), we can simply write
${\bm{\kappa}}={\bm{\kappa}}(\mbox{\boldmath{$\varepsilon$}}({\bm{u}}),p).$
Furthermore, after a backward Euler semi-discretisation in time with a constant time step and rescaling appropriately, we only consider the type of equations needed to solve at each time step and therefore we will concentrate on the form
$c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})-\mathop{\rm div}\nolimits({\bm{\kappa}}(\mbox{\boldmath{$\varepsilon$}}({\bm{u}}),p)\nabla p)=g\qquad\text{in $\Omega$}.$ (1.5)
@@
${\bm{\kappa}}(\mbox{\boldmath{$\varepsilon$}}({\bm{u}}),p)=\frac{k_{0}}{\mu_{f}}\mathbf{I}+\frac{k_{1}}{\mu_{f}}\exp(k_{2}(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})))\mathbf{I},\quad{\bm{\kappa}}(\mbox{\boldmath{$\varepsilon$}}({\bm{u}}),p)=\frac{k_{0}}{\mu_{f}}\mathbf{I}+\frac{k_{1}(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}}))^{3}}{\mu_{f}(1-(c_{0}p+\alpha\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})))^{2}}\mathbf{I},$ (1.6)
where $\mu_{f}$ denotes the viscosity of the interstitial fluid and $k_{0},k_{1},k_{2}$ are model constants. We note that in the case of incompressible constituents one has $c_{0}=0$ and $\alpha=1$, indicating that permeability depends only on the dilation $\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})=\mathop{\rm div}\nolimits{\bm{u}}$ (see, e.g., [11]). We also note that even in such a scenario (of incompressible phases) the overall mixture is not necessarily incompressible itself. More precise assumptions on the behaviour of the permeability are postponed to Section 2.2. Next we note that from (1.2) we can obtain
$\mathop{\rm tr}\bm{\sigma}=(d\lambda+2\mu)\mathop{\rm div}\nolimits{\bm{u}}-d\alpha p{\quad\hbox{and}\quad}\mathcal{C}^{-1}\bm{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbf{I}=\mbox{\boldmath{$\varepsilon$}}({\bm{u}})\qquad\text{in $\Omega$}.$ (1.7)
@@
$\mathop{\rm tr}\mbox{\boldmath{$\varepsilon$}}({\bm{u}})=\dfrac{1}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}+\dfrac{d\alpha}{d\lambda+2\mu}p,$
@@
$\begin{array}[]{cc}\displaystyle{\bm{\kappa}}(\bm{\sigma},p)=\frac{k_{0}}{\mu_{f}}\mathbf{I}+\frac{k_{1}}{\mu_{f}}\exp\Big{(}\dfrac{k_{2}}{d\lambda+2\mu}\big{(}(c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\bm{\sigma}\big{)}\Big{)}\mathbf{I},\\[4.30554pt] \displaystyle{\bm{\kappa}}(\bm{\sigma},p)=\frac{k_{0}}{\mu_{f}}\mathbf{I}+\frac{k_{1}\big{(}(c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\bm{\sigma}\big{)}^{3}}{(d\lambda+2\mu)\mu_{f}\big{(}d\lambda+2\mu-((c_{0}(d\lambda+2\mu)+d\alpha^{2})\,p+\alpha\mathop{\rm tr}\bm{\sigma})\big{)}^{2}}\mathbf{I},\end{array}$ (1.8)
and emphasize that the permeability tensor is the same constitutive law on either (1.6) or (1.8), so, making abuse of notation we have
${\bm{\kappa}}={\bm{\kappa}}(\zeta)={\bm{\kappa}}(\mbox{\boldmath{$\varepsilon$}}({\bm{u}}),p)={\bm{\kappa}}(\bm{\sigma},p).$
@@
$\mathcal{C}^{-1}\bm{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbf{I}=\bm{\nabla}{\bm{u}}-\bm{\rho}\qquad\text{in $\Omega$}.$ (1.9)
Finally, we introduce the discharge flux ${\bm{\varphi}}$ as an unknown defined by the constitutive relation
${\bm{\kappa}}(\bm{\sigma},p)^{-1}{\bm{\varphi}}=\nabla p,$ (1.10)
@@
$c_{0}p+\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}p-\mathop{\rm div}\nolimits{\bm{\varphi}}=g\qquad\text{in $\Omega$}.$ (1.11)
To close the system, we consider mixed boundary conditions for a given ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1/2}(\Gamma_{\mathrm{D}})$ and $p_{\mathrm{D}}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{D}})$:
${\bm{u}}={\bm{u}}_{\mathrm{D}}{\quad\hbox{and}\quad}p=p_{\mathrm{D}}{\quad\hbox{on}\quad}\Gamma_{\mathrm{D}},\qquad{\bm{\varphi}}\cdot\mbox{\boldmath{$n$}}=0{\quad\hbox{and}\quad}\bm{\sigma}\mbox{\boldmath{$n$}}=\bm{0}{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}.$ (1.12)